OpenStudy (anonymous):

3^3/3^6 A 1/27 B 1/9 C 9 D -27

OpenStudy (anonymous):

@Kitten_is_back

OpenStudy (anonymous):

A

OpenStudy (anonymous):

i dont want the answer

OpenStudy (er.mohd.amir):

a^n/a^m=a^(n-m) 3^-3=1/3^3=1/27

OpenStudy (anonymous):

3^3 = 3*3*3 = 27 3^6 = 3*3*3*3*3*3 = 729 27/729 = 0.03... 1/27 = 0.03... Answer = 1/27

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem.